OpenStudy (anonymous):

Find the midpoint of the segment with the given endpoints. M(–2, 1) and O(–3, 2)

OpenStudy (anonymous):

M= (y2-y1)/(x2-x1)

OpenStudy (anonymous):

still dont understand it

OpenStudy (anonymous):

Sub your points in.

OpenStudy (anonymous):

i trie that but i got -1 and 0 dont think im doing it right

OpenStudy (anonymous):

tried*

OpenStudy (anonymous):

do i do the m as the x or just the first one..

OpenStudy (anonymous):

Oops! Read your question wrong! (Sorry, my english is not that good) To find the x-value of your midpoint use this: (x1+x2)/2 To find your y-value of your midpoint use this: (y1+y2)/2

OpenStudy (anonymous):

X=-3+(-2)/2 =-5/2 Y=2+1/2 =3/2 M=(-5/2 , 3/2)